OpenStudy (barrelracing):

Solve −2x2 +3x − 9 = 0

OpenStudy (arindameducationusc):

you need x?

OpenStudy (barrelracing):

a) \[ \frac{3\pm3i \sqrt{7} }{ 4 }\]

OpenStudy (arindameducationusc):

yes, I guess so...

OpenStudy (barrelracing):

b) \[\frac{ 3\pm9i \sqrt{7} }{ 4 }\]

OpenStudy (arindameducationusc):

A is correct

OpenStudy (barrelracing):

c) \[\frac{ -3\pm3i \sqrt{7} }{ 4 }\]

OpenStudy (barrelracing):

d) \[\frac{ -3\pm9i \sqrt{7} }{ 4}\]

OpenStudy (barrelracing):

are you sure?

OpenStudy (barrelracing):

@freckles

OpenStudy (arindameducationusc):

Yes I am sure

OpenStudy (arindameducationusc):

I slept yesterday, sorry about that